# Randomly Sampled Language Reasoning Problems Reveal Limits of LLMs

### v4

@@ Abstract
Can LLMs pick up language structure from examples? Evidence in prior work seems to indicate yes, as pretrained models repeatedly demonstrate the ability to adapt to new language structures. However, this line of research typically considers languages that are present within common pretraining datasets, or otherwise share notable similarities with seen languages. In contrast, in this work we attempt to measure models’ language understanding capacity while circumventing the risk of dataset recall. We parameterize large families of language tasks recognized by deterministic finite automata (DFAs), and can thus sample novel language reasoning problems to fairly evaluate LLMs regardless of training data. We find that, even in the strikingly simple setting of 3-state DFAs, LLMs underperform unparameterized $n$-Gram models on both language recognition and synthesis tasks. These results suggest that LLMs struggle to match the ability of basic language models in recognizing and reasoning over languages that are sufficiently distinct from the ones seen at training time, underscoring the distinction between learning individual languages and possessing a general theory of language.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: We sample randomly generated languages to test LLMs by sampling deterministic finite automata (DFAs). (a) The DFA shown here, modeling the sum modulo 3 operation (with abc representing 0, 1, and 2 respectively), can be used to accept or reject strings from a 3-character alphabet. Accepted strings belong to the grammar, and rejected strings do not. We evaluate models on their ability to (b) act as a transducer, recognizing strings that belong to the DFA-defined grammar, and (c) generate new strings following the grammar.
Contemporary LLMs have proven themselves to be highly sophisticated natural language completion models that demonstrate many properties of reasoning engines. This has prompted questions surrounding the true intelligence of these models, with some arguing that they possess inherent language learning capabilities (Millière, 2024). In this paper, we explore the question of whether LLMs have the reasoning capacity to understand the structure of a new language. Specifically, we are interested in problems where a model is given a small set of examples from a language and either generates a new sample or determines whether a new sample is from the language or not.
Some work suggests that LLMs broadly understand language structure because they are able to produce syntactically correct samples from languages they have not been trained on (Athiwaratkun et al., 2022), although LLMs’ performance on low-resource languages tends to be lower than their ability on higher resource languages (Bogin et al., 2023). On the other hand, some critics of LLMs argue that LLMs cannot possess an understanding of language structure as they have learned from data rather than possessing a priori universal grammar (Chomsky et al., 2023). We do not view learning from data as a fundamental limitation, but we are concerned with the possibility that a language model might only be able to understand linguistic structures similar to those it has seen in training data. To distinguish between these possibilities, we wish to evaluate LLMs on a benchmark that considers wholly novel languages, eliminating the possibility of dataset leakage.
Ideally, models would be tested on a set of language reasoning problems disjoint from data seen during training and validation. However, as training datasets for LLMs are generally closed and incredibly vast, human-generated problems in natural language are likely to at least partially overlap in syntax or concept with content LLMs have already seen, making probing a model for its ability to reason about the structure of natural languages nearly impossible. Additionally, the problem of determining whether two tasks are semantically identical is itself a nontrivial one. Therefore, ensuring that even an entirely novel invented problem is not a variation on a theme is intractable.
@@
To circumvent this problem, we propose the following general approach: first we define a large, exhaustive, and parsimoniously-defined space of languages that represents all languages of a certain difficulty level. Then, we sample random languages from this space. By sampling randomly, we can guarantee no bias towards canonical languages that might share structure with common ones in the training dataset. In this work, we use languages recognized by 3-state DFAs as these are the lowest nontrivial difficulty level, but this technique can be generalized to produce benchmarks of any difficulty level.
Experimental results using this approach suggest that contemporary LLMs possess less sophisticated language pattern recognition abilities than expected; underperforming basic, parameter-free n-gram language models on even the simplest languages. These results, combined with LLMs’ impressive results on a variety of specific tasks, suggest that LLMs function as ensemble models over language tasks they have seen in their dataset, but do not possess the ability to generalize to entirely novel language reasoning tasks.
@@
    We introduce a benchmark for LLM language reasoning evaluation, disjoint from natural language web data.
@@ 2.2 LLM reasoning evaluation
LLM reasoning abilities are often tested on natural language benchmarks and commonly seen problems like arithmetic (Cobbe et al., 2021; Amini et al., 2019; Hendrycks et al., 2021), commonsense reasoning (Bhargava & Ng, 2022), and other, sometimes generative, tasks (Lake & Baroni, 2018; Pasupat & Liang, 2015; Lin et al., 2019) and task collections (Srivastava et al., 2022). LLMs have been shown to lack sufficient reasoning capability across a range of tasks including multi-step planning and complex inference (Valmeekam et al., 2022). Fan et al. (2023) introduce an LLM reasoning benchmark on algorithmic problems through NP-hard complexity, and Hazra et al. (2024) show that LLMs struggle to complete simple 3SAT problems. Patel et al. (2021) demonstrate that much of LLM mathematical reasoning can be explained by shallow heuristics, and Razeghi et al. (2022) similarly find that term frequency in training data impacts models’ in-context learning ability. In comparison to these, we explore the distinction described by Patel et al. (2021), but push both language simplicity and language unfamiliarity to their limits, by exploring simple languages recognized by randomly sampled DFAs. This enables us to evaluate the ability of LLMs to reason about language.
2.3 Language Understanding and LLMs
LLMs can be quite adept at generating programs in general-purpose programming languages (Xu et al., 2022a). In contrast, adapting models to understand domain-specific languages (Mernik et al., 2005) introduces unique problems such as navigating idiosyncratic syntax and semantics, and leveraging sparse collections of sample language data. To address these challenges, researchers have considered how well general-purpose LLMs can use language reasoning skills to quickly understand rare or unseen DSLs with only a small set of exemplars (Joel et al., 2024). While most work in this vein focuses on semantic parsing for downstream applications (Lin et al., 2023), selecting exemplars (Zhao et al., 2021), and improving DSL recognition by leveraging more common languages (Bogin et al., 2023), experiments show strong baseline performance for LLM DSL recognition and parsing out-of-the-box (Wang et al., 2024), indicating that LLMs may possess emergent language reasoning abilities.
@@ 2.4 Training transformers on Formal Languages
Vafa et al. (2025) frame world modeling as a latent DFA identification task, finding that transformers trained on DFA traces (of massive DFAs representing board games and city maps) do not reconstruct the underlying DFA. Akyürek et al. (2024) find that transformers trained on 4-12 state DFA transducer traces more effectively learn to in-context-learn regular languages than RNNs. Here, we evaluate much larger and more capable LLMs on much simpler 3-state DFAs. However, our goal is evaluating the abilities of foundation models (trained on a broad corpus rather than a tailored one), and is thus not strictly easier.
@@ 3.1 DFAs and Regular Languages
The original Chomsky Hierarchy (Chomsky, 1959) separates language into four types (Figure 2). We focus on the task of understanding Type 3 languages, the simplest form of language in the hierarchy, that are recognized by a Deterministic Finite Automaton (DFA). Examples of languages recognized by DFAs include simple ones like binary strings with an even number of ones, and even such examples as numbers in base 10 divisible by 7. Type 3 languages are also known as regular languages, which are recognized by regular expressions.
@@
3.2 Sequence Completion Task
We first pose a sequence completion task, in which models must complete a sequence in a given DFA’s language. In practice, most language data that models encounter will be in roughly this format, with several example sequences in a given language followed by a distinct prefix that needs to be completed via next token prediction.
To generate test cases for this task given a DFA, we first (1) sample 30 example sequences of length 10 that this DFA accepts, and then (2) sample a distinct prefix of length 5 that is not a prefix of any of our 30 example sequences, with the property that there exists some length-$\leq 5$ completion of this prefix that the DFA would accept. The task is to find a completion (not necessarily the same completion found in sampling) of this prefix of between 1 and 5 characters such that the DFA accepts the full sequence. For details on sampling, see Appendix A.2.
@@ 3.2 Sequence Completion Task
3.3 Transducer Task
While the sequence completion task is the natural one that comes to mind as a basic language task, it has a difficulty-gap problem. Specifically, the issue is that many DFAs, including the one shown in Figure 1, recognize languages that are particularly difficult to identify based on a set of examples, unless you build some kind of world model. Other DFAs end up being trivial to generate a completion for by analyzing common suffixes.222 The difficulty gap exists because a set of recognized sequences of length 10 gives no direct insight into intermediate states between the first and tenth token. As such, to be able to utilize this information for languages like the one in Figure 1 where there are no “resets” (sequences of symbols that necessarily lead to a particular state), a model must be capable of hollistically evaluating the entire sequence, probably requiring a world model. Many other DFAs contain these resets, but do so in such a way that makes it possible to e.g., recognize that all sequences that end in a are in the language, making the problem trivial. To provide a more direct evaluation of non-world-modeling-based pattern recognition, we explore the Transducer task.
In this task, an input sequence is annotated with an output at each token, the final output is masked, and the masked output is predicted by a language model. We call this a transducer task, as the DFA converts a sequence of inputs into a sequence of outputs. E.g., given the language does the string have an even number of ’a’ tokens and the input abcabcaabbccaa, the annotated string (all that is provided to the model) is a0b0c0a1b1c1a0a1b1b1c1c1a0a and the output to predict is 1. For each problem instance, we provide 30 symbols, and for the first 29, the corresponding transducer output.
This task is significantly more transparent than the sequence completion task as the model has access to intermediate outputs, an (imperfect) proxy for intermediate state. n
@@ 3.4 Baselines
l To contextualize LLM accuracies, waseline models with varying degrees of sophistication.
@@
For the Sequence Completion task, we have three kinds of baseline.
@@
  • Common-SuffixS baseline: find the completion $s$ of length between 1 and 5 that maximizes ((((# of occurrences as a ${}\times|s|)$. This baseline does not take the distinct prefix into account, and instead tries to find a universal completion that will always end in an accept state for this language.
  • $n$-GramS baseline: we take the last $n-1$ characters of the distinct prefix and search to see if they appear in any of the example sequences at a position where the sequence following is an appropriate length to be a completion (at least 1 but at most 5). We then take a plurality vote among the completions and return this, breaking ties arbitrarily. If there are no matches, we return the result of $(n-1)$-GramS. Technically these cover more than $n$ characters, since the completion is often $>1$ character long; for simplicity, however, we keep the naming consistent with the Transducer baselines.
@@
Note that these baselines are entirely unparameterized and operate identically regardless of the underlying DFA. This makes them direct comparisons to using LLMs in in-context-learning. We do not consider BruteForceS to be a reasonable comparison due to its computational complexity, and instead consider it an upper bound on performance on this particular task.
@@
  • $n$-GramT baseline: take the $n-1$ symbols ending at the end of the concatenated transducer sequence (e.g., for $n=5$ and the above example, this would be 1a0a). If that sequence does not appear elsewhere in the sequence, return the result of the $(n-1)$-GramT baseline. Otherwise, take the token that appears immediately after each occurrence. If there is a majority, return that, otherwise return the last example.
@@ 4 Experiments
We evaluated the open-source models Llama 3-8B, Llama 3-70B (AI@Meta, 2023), and Llama 3-8B-Instruct (AI@Meta, 2024), Mistral Nemo Minitron 8B (NVIDIA, 2024), Mistral Nemo Base 2407 (Mistral AI, 2024b) and Mistral Nemo Instruct 2407 (Mistral AI, 2024c), Gemma 7B (Google, 2024), and Falcon 7B (Almazrouei et al., 2023).
@@
We also evaluated the open-source code models StarCoder2-15B (Lozhkov et al., 2024), Codestral-22B-v0.1 (Mistral AI, 2024a), Deepseek Coder 33B Instruct (Deepseek, 2024), Qwen2.5-Coder-7B, Qwen2.5-Coder-7B-Instruct, and Qwen2.5-Coder-32B-Instruct (Hui et al., 2024).
@@
Finally, we evaluated the proprietary models GPT-3.5-turbo-instruct, GPT-3.5 Chat (turbo-0125) (OpenAI, 2024a), GPT-4o-mini (2024-07-18), GPT 4o (2024-05-13) (OpenAI, 2024b), o1-preview (2024-09-12) (OpenAI et al., 2024), and Claude 3.5 Sonnet (Anthropic, 2024).
For each open source model, we used a local VLLM (Kwon et al., 2023) server for evaluation and always evaluated on 1000 distinct DFAs. For GPT-4o and Claude, we evaluated on 30 DFAs due to computation costs. For o1-preview we evaluated on only 10 DFAs, and only on the Transducer task (which we felt was a better fit for a reasoning model). For gpt-3.5 and gpt-4o-mini, we evaluated on 100 DFAs. All models were evaluated with temperature 0, except o1-preview333o1-preview does not allow setting a non-default temperature..
For both tasks, we consider four prompting formats:
@@
  • Basic provides no context, presenting the problem as a generic sequence generation or next-token prediction task, where output is provided immediately following the input, with no space to think.
  • More-Expl explains that the strings are generated from a DFA with 3 states, but is otherwise identical to Basic. This remains a sequence generation/next token prediction task.
  • COT provides the same information as More-Expl and additionally invokes chain-of-thought reasoning to help the model reason over the task. Here, the model is given space to reason before providing a tagged answer.
@@
  • Red-Green casts the tasks as independent word problems that describe the underlying grammar structure without relying on world knowledge about DFAs and regular languages. It describes an N-state DFA as a house with N rooms, each of which has 3 portals that deterministically go to other rooms (or back to the same room), where the walls of each room are red or green (mirroring transducer output symbols 0 and 1). Similarly to COT, the model is given space to show work before providing a tagged answer.
@@
We produce versions of each of these prompts for each task, denoting these with a subscript S for sequence completion prompts and T for transducer prompts. Full listings of these prompts can be found in Appendix F. While no finite set of prompts will be fully sufficient to capture all possible model behavior, we believe our set of prompts captures common prompting strategies.
@@
RandomS/NullT 53.3 (51.7–54.7) 26 68.9 (68.2–69.6) 26
Open Source Completion
llama3-8B 8.0B 73.8 (72.4–75.1) 18 87.5 (86.9–88.0) 14
llama3-70B 70.6B 71.4 (70.0–72.7) 23 87.7 (87.2–88.3) 12
llama3.1-8B-Instruct 8.0B 75.3 (74.0–76.6) 16 85.9 (85.3–86.5) 18
mistral-nemo-minitron-8B 8.4B 78.7 (77.5–79.8) 12 88.6 (88.0–89.1) 5
mistral-nemo-base-12B 12.2B 75.5 (74.3–76.6) 15 87.9 (87.4–88.4) 10
mistral-nemo-instruct-12B 12.2B 72.2 (70.9–73.4) 22 88.0 (87.5–88.5) 8
gemma-7b 8.5B 72.6 (71.3–73.7) 20 82.1 (81.4–82.7) 22
falcon-7b 7.2B 69.0 (67.6–70.2) 24 84.9 (84.3–85.5) 20
Open Source Code
starcoder2-15b 16.0B 73.5 (72.0–74.7) 19 87.7 (85.8–89.5) 13
codestral-22B 22.2B 78.0 (76.8–79.1) 13 86.6 (86.0–87.1) 17
deepseek-coder-33b-instruct 33.3B 76.7 (75.3–77.8) 14 85.6 (85.0–86.2) 19
qwen-2.5-coder-7B 7.6B 79.5 (78.4–80.5) 9 88.2 (87.6–88.7) 7
qwen-2.5-coder-instruct-7B 7.6B 79.5 (78.3–80.5) 10 88.3 (87.8–88.8) 6
qwen-2.5-coder-instruct-32B 32.8B 79.2 (78.0–80.3) 11 87.9 (87.4–88.4) 9

### v6

# Randomly Sampled Language Reasoning Problems Elucidate Limitations of In-Context Learning

@@ Abstract
While LLMs have revolutionized the field of machine learning due to their high performance on a strikingly wide range of problems, they are also known to hallucinate false answers and underperform on less canonical versions of the same tasks. There are several emerging theories of LLM performance, among them that LLMs lack world modeling ability, that they have an undesirable bias towards an autoregressive prior, and that they struggle on more novel problems. The existing literature on LLM input novelty has focused on tasks of relatively high complexity, studying perturbations of canonical but complex problems. In this paper, we attempt to minimize complexity in order to isolate novelty as a factor in LLM underperformance and investigate the power of in-context-learning. To this end, we consider an extremely simple domain: next token prediction on simple language tasks. The twist is that these language tasks are wholly unseen, as they are randomly drawn from a large, parsimoniously defined set of languages arising from simple grammar rules. This experimental setup allows us to evaluate ICL independently of models’ parametric knowledge. We find that LLMs uniformly underperform n-gram models on this task, both when used as next token predictors and in chain-of-thought.
@@ 1 Introduction
Figure 1: We sample randomly generated languages to test LLMs by sampling deterministic finite automata (DFAs). (a) The DFA shown here, modeling the sum modulo 3 operation (with abc representing 0, 1, and 2 respectively), can be used to accept or reject strings from a 3-character alphabet. Accepted strings belong to the grammar; rejected strings do not. We evaluate models on their ability to (b) act as a transducer, recognizing strings that belong to the grammar, and (c) generate new strings following the grammar.
One of the surprising capabilities of contemporary LLMs is their ability to perform in-context-learning (ICL), in which they learn by demonstration from provided input/output examplars to produce an appropriate output for a new but similarly structured input. This capability allows LLMs to generalize to tasks beyond interpolations of their training corpus, resulting in remarkably strong generalization capabilities. It is challenging to isolate the effects of ICL, as, given the large datasets LLMs are trained on, it is difficult to distinguish a model detecting novel patterns in examples from it being guided towards knowledge it gained during training. Akyürek et al. (2022) demonstrate that, in specially trained transformers, ICL is “true learning”. However, it remains unclear to what degree the performance of foundation-model LLMs can be attributed to strong ICL capabilities.
The task best suited to isolating LLM ICL performance should have the following properties: (1) It should be a language completion task within the expressive power of an LLM. LLMs are capable of many tasks, but they are primarily models of language, and as such, language tasks are the most fair evaluations. (2) It should not require sophisticated world modeling to solve. This helps us eliminate a possible source of underperformance distinct from ICL capability. (3) It should be selected at random in an unbiased manner to reduce the effect of bias from the training corpus.
@@
To satisfy these properties, we propose the following general approach. First we define a large, exhaustive, and parsimoniously-defined space of languages that represents all languages of a certain difficulty level. Then, we sample random languages from this space. By sampling randomly, we can guarantee no bias towards canonical languages that might share structure with common ones in the training dataset. In this work, we use languages recognized by 3-state DFAs as these are the lowest nontrivial difficulty level.  [^1] This technique can be generalized to produce benchmarks of any difficulty level. For larger numbers of states, we would be able to guarantee that the majority of the exponentially large number of corresponding languages do not lie in the training dataset by a pigeonhole argument; unfortunately this does not apply to the relatively small set of 3-state DFAs (there are only 78786). However, they still represent an set of tasks of a particular difficulty level not biased towards the canonical. Finally, to ensure we are not measuring world modeling performance, we compare to $n$-Gram baselines that are not capable of anything other than matching clusters of tokens.
Our results demonstrate that even for very simple language induction tasks that don’t rely on world modeling or background knowledge, LLM ICL still underperforms simple language models when dealing with randomly sampled and likely unfamiliar problem instances. These results suggest that while LLMs can pick up some learning signal from examples in prompts, this in-context learning is not competitive with even very primitive forms of learning, and suggests that LLMs do not posses the ability to generalize to entirely novel language reasoning tasks.
@@
    We introduce a benchmark for LLM ICL language reasoning evaluation, consisting of novel tasks.
@@ 2 Related Work
LLMs are known to fail in many cases, with some suggesting that these failures are due to lack of a world model (Valmeekam et al., 2022) or “embers” of autoregression polluting non-autoregressive task performance (McCoy et al., 2023). Another theory is that of task novelty; that is, LLMs perform worse on tasks more dissimilar from those seen during training.
2.1 Language Understanding and LLMs
LLMs can be quite adept at generating programs in general-purpose programming languages (Xu et al., 2022a). In contrast, adapting models to understand domain-specific languages (Mernik et al., 2005) introduces unique problems such as navigating idiosyncratic syntax and semantics and leveraging sparse sample language data. To address these challenges, researchers have considered how well general-purpose LLMs can use language reasoning skills to quickly understand rare or unseen DSLs with only a small set of exemplars (Joel et al., 2024). While most work in this vein focuses on semantic parsing for downstream applications (Lin et al., 2023), selecting exemplars (Zhao et al., 2021), and improving DSL recognition by leveraging more common languages (Bogin et al., 2023), experiments show strong baseline performance for LLM DSL recognition and parsing out-of-the-box (Wang et al., 2024). Some have suggested that indicate that LLMs may possess emergent language reasoning abilities (Millière, 2024).
@@ 2.3 LLM reasoning evaluation
LLM reasoning abilities are often tested on natural language benchmarks and commonly seen problems like arithmetic (Cobbe et al., 2021; Amini et al., 2019; Hendrycks et al., 2021), commonsense reasoning (Bhargava & Ng, 2022), and other, sometimes generative, tasks (Lake & Baroni, 2018; Pasupat & Liang, 2015; Lin et al., 2019) and task collections (Srivastava et al., 2022). LLMs have been shown to lack sufficient reasoning capability across a range of tasks including multi-step planning and complex inference (Valmeekam et al., 2022). Fan et al. (2023) introduce an LLM reasoning benchmark on algorithmic problems through NP-hard complexity, and Hazra et al. (2024) show that LLMs struggle to complete simple 3SAT problems. Patel et al. (2021) demonstrate that much of LLM mathematical reasoning can be explained by shallow heuristics, Razeghi et al. (2022) similarly find that term frequency in training data impacts models’ in-context learning ability, and (McCoy et al., 2023) theorizes that “embers” of autoregression are polluting non-autoregressive task performance.
The effect of novelty on performance has also been explored in prior work, generally via investigating perturbations of existing tasks, taking existing problems (that are often inherently quite complex, and are only easy because they are well known, e.g., addition of numbers or logical reasoning over natural language) and changing one small aspect of the problem (Wu et al., 2024; Saparov et al., 2023). We also study task novelty but push both language simplicity and language unfamiliarity to their limits, by exploring simple languages recognized by randomly sampled DFAs. This enables us to best isolate the power of ICL in the language domain, where LLMs should perform best.
@@ 2.4 Training transformers on Formal Languages
A key assumption behind this work is that the tasks we are using to evaluate LLMs are solvable by LLMs. Vafa et al. (2025) frame world modeling (a statistical model inferring the true underlying causal graph behind the data being observed) as a latent DFA identification task, finding that transformers trained on DFA traces (of massive DFAs representing board games and city maps) do not reconstruct the underlying DFA. Other work also trains language models on formal languages (Butoi et al., 2024; Bhattamishra et al., 2023; Valvoda et al., 2022) and probabilistic formal languages (Borenstein et al., 2024). Akyürek et al. (2024) find that transformers trained on 4-12 state DFA transducer traces more effectively learn to in-context-learn regular languages than RNNs or $n$-Gram models. Therefore, in this work, where we evaluate much larger LLMs on much simpler 3-state DFAs, we can be confident that underperformance relative to $n$-Grams is not linked to inherent transformer limitations and must be instead related somehow to specific properties of foundation models.
@@ 3.1 DFAs and Regular Languages
The original Chomsky Hierarchy (Chomsky, 1959) separates language into four types (Figure 2). We focus on the task of understanding Type 3 languages, the simplest form of language in the hierarchy, that are recognized by a Deterministic Finite Automata (DFAs) whose outputs are boolean ({0, 1}). Examples of languages recognized by DFAs include simple ones like binary strings with an even number of ones, and even such examples as numbers in base 10 divisible by 7. Type 3 languages are also known as regular languages, which are recognized by regular expressions.
@@
3.2 Language Reasoning Tasks
We define a language reasoning task as a task corresponding to some latent language $\mathcal{L}$, where a set of positive/negative examples is provided to a model, with the goal being either classification of a new string in this language, or completion of an existing string to place it in the language.
3.2.1 Sequence Completion Task
We first pose a sequence completion task, in which models must complete a sequence in a given DFA’s language. This mirrors how foundation models are trained using masked language modeling, where data is presented in this format, with several example sequences in a given language followed by a distinct prefix that needs to be completed.
To generate test cases for this task given a DFA, we (1) sample 30 example sequences of length 10 that this DFA accepts, and then (2) sample a distinct prefix of length 5 that is not a prefix of any of our 30 example sequences, with the property that there exists some length-$\leq 5$ completion of this prefix that the DFA would accept. The task is to find a completion (not necessarily the same completion found in sampling) of this prefix of between 1 and 5 characters such that the DFA accepts the full sequence. For details on sampling, see Appendix A.2.
@@ 3.2.1 Sequence Completion Task
3.2.2 Transducer Task
While the sequence completion task is the natural one that comes to mind as a basic language task, it has a difficulty-gap problem. Specifically, many DFAs, including the one shown in Figure 1, recognize languages that are particularly difficult to identify based on a set of examples, unless you build some kind of world model.333 The difficulty gap exists because a set of recognized sequences of length 10 gives no direct insight into intermediate states between the first and tenth token. As such, to be able to utilize this information for languages like the one in Figure 1 where there are no “resets” (sequences of symbols that necessarily lead to a particular state), a model must be capable of hollistically evaluating the entire sequence, probably requiring a world model. Many other DFAs contain these resets, but do so in such a way that makes it possible to e.g., recognize that all sequences that end in a are in the language, making the problem trivial. This is problematic as we would like to be able to assess the performance of language models at pattern recognition, independent of their world modeling abilities. To assess pattern recognition, we explore the Transducer task.
In this task, an input sequence is annotated with an output at each token, the final output is masked, and the masked output is predicted by a language model. E.g., given the language even number of ’a’ tokens and the input abcabcaabbccaa, the annotated string (all that is provided to the model) is a0b0c0a1b1c1a0a1b1b1c1c1a0a and the output to predict is 1. For each problem instance, we provide 30 symbols, and for the first 29, the corresponding transducer output.
This task is significantly more transparent than the sequence completion task as the model has access to intermediate outputs, an (imperfect) proxy for intermediate state.
@@ 3.3 Baselines
To contextualize LLM accuracies, we provide several baselines with varying degrees of sophistication.
@@
For the Sequence Completion task, we have four baselines.
@@
[…]
  • $n$-GramS baseline: we take the last n1n-1 characters of the distinct prefix and search to see if they appear in any of the example sequences at a position where the sequence following is an appropriate length to be a completion (at least 1 but at most 5). We then take a plurality vote among the completions and return this, breaking ties arbitrarily. If there are no matches, we return the result of (n1)(n-1)-GramS. Technically these cover more than $n$ characters, since the completion is often >1>1 character long; for simplicity, however, we keep the naming consistent with the Transducer baselines. Despite the similarity between an $n$-Gram and a DFA in terms of token-to-token transitions, $n$-Grams do not have access to DFA hidden state and thus cannot solve arbitrary DFA language problems, regardless of $n$.
@@
Note that these baselines are entirely unparameterized and operate identically regardless of the underlying DFA. This makes them direct comparisons to using LLMs in in-context-learning [^4] One thing to note is that LLMs are required to determine that they are performing next token prediction on a particular string from a natural language description such as “You are a sequence completion model…,” while $n$-Gram models are programmed to do so. However, we believe all LLMs we evaluate are sophisticated enough to accomplish this without issue.. We do not consider BruteForceS to be a reasonable comparison due to its computational complexity, and instead consider it an upper bound on performance on this particular task. We choose $n$-Gram baselines as they are are unambiguously representable by transformers (Svete & Cotterell, 2024), so a transformer model should be able to match their performance.
@@
  • $n$-GramT baseline: take the n1n-1 symbols ending at the end of the concatenated transducer sequence (e.g., for $n=5$ and the above example, this would be 1a0a). If that sequence does not appear elsewhere in the sequence, return the result of the (n1)(n-1)-GramT baseline. Otherwise, take the token that appears immediately after each occurrence. If there is a majority, return that, otherwise return the last example.
@@ 4 Experiments
We evaluated the open-weight models Llama 3-8B, Llama 3-70B (AI@Meta, 2023), Llama 3.1-8B (AI@Meta, 2024b), Llama 3.1-8B-Instruct (AI@Meta, 2024c), Llama 3.1-70B (AI@Meta, 2024a), Mistral Nemo Minitron 8B (NVIDIA, 2024), Mistral Nemo Base 2407 (Mistral AI, 2024b) and Mistral Nemo Instruct 2407 (Mistral AI, 2024c), Gemma 7B (Google, 2024), Falcon 7B (Almazrouei et al., 2023), Qwen 2.5-7B and Qwen 2.5-32B Team (2024).
@@
We also evaluated the open-weight code models StarCoder2-15B (Lozhkov et al., 2024), Codestral-22B-v0.1 (Mistral AI, 2024a), Deepseek Coder 33B Instruct (Deepseek, 2024), Qwen2.5-Coder-7B, Qwen2.5-Coder-7B-Instruct, and Qwen2.5-Coder-32B-Instruct (Hui et al., 2024).
@@
Finally, we evaluated the proprietary models Claude 3.5 Sonnet (Anthropic, 2024), GPT-3.5-turbo-instruct, GPT-3.5 Chat (turbo-0125) (OpenAI, 2024a), GPT-4o-mini (2024-07-18), GPT 4o (2024-05-13) (OpenAI, 2024b), o3-mini (2025-01-31) (OpenAI, 2025b), and gpt-5 (2025-08-07) (OpenAI, 2025a).
@@
For both tasks, we consider two main prompting formats. Basic provides no context, presenting the problem as a generic sequence generation or next-token prediction task, where output is provided immediately following the input, with no space to think. Basic-COT provides the same prompt but asks the model to think step by step and provide an answer. These prompts test ICL, presenting the task in an unstructured manner and requiring the model to learn the problem structure via induction. Our main results are the maximum over these two prompting strategies.
@@
We also provide three “control” prompting formats where information on the problem structure is provided. More-Expl explains that the strings are generated from a simple grammar, but is otherwise identical to Basic. This remains a sequence generation/next token prediction task. DFA-COT provides the full structure of the latent language, stating that it is a 3-state DFA, and additionally invokes chain-of-thought reasoning to help the model reason over the task. Red-Green casts the tasks as independent word problems that describe the underlying grammar structure without relying on world knowledge about DFAs and regular languages. It describes an N-state DFA as a house with N rooms, each of which has 3 portals that deterministically go to other rooms (or back to the same room), where the walls of each room are red or green (mirroring transducer output symbols 0 and 1). Similarly to DFA-COT, the model is given space to show work before providing a tagged answer.
@@
We produce versions of each of these prompts for each task, denoting these with a subscript S for sequence completion prompts and T for transducer prompts. Full listings of these prompts can be found in Appendix I. While no finite set of prompts will be fully sufficient to capture all possible model behavior, we believe maximizing over both Basic prompts allows both models that perform best at next-token-prediction and those that perform better in chain-of-thought reasoning to do their best.
For each open weight model, we used a local VLLM (Kwon et al., 2023) server for evaluation and always evaluated on 1000 distinct DFAs. For GPT-4o and Claude, o3-mini, and gpt-5, we evaluated on 30 DFAs due to computation costs. (Due to greater interest in o3-mini’s performance on Red-GreenT, we used 100 to get a more precise estimate). For gpt-3.5 and gpt-4o-mini, we evaluated on 100 DFAs. All models were evaluated with temperature 0, except reasoning models o3-mini and gpt-5, which do not support a custom temperature.
@@
RandomS/NullT 53.3 (51.7–54.7) 32 68.9 (68.2–69.6) 31
Open Weight Completion
llama3-8B 8.0B 73.8 (72.4–75.1) 22 87.5 (86.9–88.0) 18
llama3-70B 70.6B 71.4 (70.0–72.7) 29 87.7 (87.2–88.3) 15
llama3.1-8B-Instruct 8.0B 75.3 (74.0–76.6) 19 85.9 (85.3–86.5) 22
llama3.1-8B 8.0B 75.2 (73.8–76.3) 20 88.0 (87.5–88.6) 10
llama3.1-70B 70.0B 71.8 (70.4–73.1) 28 87.7 (87.2–88.2) 17
qwen-2.5-7B 7.6B 73.5 (72.1–74.8) 24 88.7 (88.2–89.2) 5
qwen-2.5-32B 32.5B 76.8 (75.5–78.0) 16 88.3 (87.8–88.8) 7
mistral-nemo-minitron-8B 8.4B 78.7 (77.5–79.8) 13 88.6 (88.0–89.1) 6
mistral-nemo-base-12B 12.2B 75.5 (74.3–76.6) 18 87.9 (87.4–88.4) 13
mistral-nemo-instruct-12B 12.2B 72.2 (70.9–73.4) 27 88.0 (87.5–88.5) 11
gemma-7b 8.5B 72.6 (71.3–73.7) 25 82.1 (81.4–82.7) 27
falcon-7b 7.2B 69.0 (67.6–70.2) 30 84.9 (84.3–85.5) 24
Open Weight Code
starcoder2-15b 16.0B 73.5 (72.0–74.7) 23 87.7 (85.8–89.5) 16
codestral-22B 22.2B 78.0 (76.8–79.1) 14 86.6 (86.0–87.1) 21
deepseek-coder-33b-instruct 33.3B 76.7 (75.3–77.8) 17 85.6 (85.0–86.2) 23
qwen-2.5-coder-7B 7.6B 79.5 (78.4–80.5) 10 88.2 (87.6–88.7) 9
qwen-2.5-coder-instruct-7B 7.6B 79.5 (78.3–80.5) 11 88.3 (87.8–88.8) 8
qwen-2.5-coder-instruct-32B 32.8B 79.2 (78.0–80.3) 12 87.9 (87.4–88.4) 12